# Range Image-Based Implicit Neural Compression for LiDAR Point Clouds

### v1

@@ IV-C1 Impact of Patch-wise INR Architecture
The proposed depth INR exploits the patch-wise architecture, whereas the pixel-wise architecture can be used for the depth INR. Fig. 6 (a) shows the CD of the proposed patch-wise INR and pixel-wise INR architectures as a function of bitrates under the different sequences of KITTI’s LiDAR point cloud. We can see that the patch-wise depth INR achieves better CD than the pixel-wise architecture at large bitrate regimes in every LiDAR sequence. Specifically, BD-CD between the patch-wise and pixel-wise architectures is 0.032, 0.017, and 0.011 in the sequences 00-00, 00-25, and 00-50, respectively. The effects on the visual quality are shown in Figs. 4 (k), (l), (w), and (x), respectively.
Fig. 6 (b) shows the CD performance of INR-based image compression methods as a function of the learning epochs. Our patch-wise architecture boosts the convergence speed by up to 13$\times$ compared to the pixel-wise architecture, and the fast convergence results in a short encoding delay.
@@ IV-C2 Impact of Model Compression
IV-C3 Impact of Patch Size
Figure 7: Chamfer distance under the different patch sizes. Here, LiDAR sequence is 00-00.
Refer to caption (a) Seq:00-00 Refer to caption (b) $N_{p}$: 2
Refer to caption (c) $N_{p}$: 4 Refer to caption (d) $N_{p}$: 8
Refer to caption (e) $N_{p}$: 16 Refer to caption (f) $N_{p}$: 32
Refer to caption (g) $N_{p}$: 64
Figure 8: Snapshots of the reconstructed LiDAR point clouds in proposed methods under the different patch sizes $N_{p}\times N_{p}$. Here, (b)-(g) show the reconstructed point clouds of sequence 00-00.
The proposed depth INR uses the patch-wise architecture, and thus the depth image is divided into patches of size $N_{p}\times N_{p}$. We set the patch size to $16\times 16$ for best performance. This section discusses an effect of the patch size on the quality of the reconstructed point cloud.
Fig. 7 shows CD performance varying $N_{p}$, and Fig. 8 demonstrates the corresponding snapshots of the reconstructed point cloud. We consider all the variants using the same model size in Fig. 6 with a sparsity of 0.0 and $N_{b}$ of 32. As mentioned, $N_{p}=16$ is the best performance. For example, in smaller patch sizes, such as ($N_{p}=2$ and $N_{p}=4$), the patches cover larger areas, resulting in a loss of fine local detail and reducing the model’s ability to capture intricate depth variations.
@@ V Conclusion and Future Work
We proposed a novel RI-based LiDAR point cloud compression method. The proposed method is designed to efficiently compress floating-point RIs using INR-based techniques and features a sophisticated architecture that combines separated learning for mask and depth images, patch-wise learning for depth images, and model compression. Experiments on the KITTI dataset show that the proposed method improves 3D reconstruction quality at low bit rates compared to existing image, point cloud, RI, and INR-based compression methods, resulting in a BD-CD improvement of up to 0.961, and the improvement also brings better performance in the downstream task of 3D object detection.
@@
The proposed method has two limitations: encoding delay and transformation loss from RIs to 3D point clouds. While existing baselines require only a few milliseconds for encoding, implicit neural compression, including the proposed method, takes from tens of minutes to several hours. In summary, the proposed method’s long encoding delay and short decoding delay make it well-suited for offline applications of LiDAR point clouds. In addition, RIs with limited spatial resolution will lead to irreversible point loss during 2D-to-3D decoding, potentially degrading the performance of downstream tasks. In future work, we will consider integrating a point cloud generator [42] to obtain a denser point cloud from the limited resolution of RIs.

### v2

@@ IV-C1 IMPACT OF PATCH-WISE INR ARCHITECTURE
The proposed depth INR exploits the patch-wise architecture, whereas the pixel-wise architecture can be used for the depth INR. Fig. 6 (a) shows the CD of the proposed patch-wise INR and pixel-wise INR architectures as a function of bitrates under the different sequences of KITTI’s LiDAR point cloud. We can see that the patch-wise depth INR achieves better CD than the pixel-wise architecture at large bitrate regimes in every LiDAR sequence. Specifically, BD-CD between the patch-wise and pixel-wise architectures is 0.047, 0.031, and 0.028 in frames 00, 25, and 50 of sequence 00, respectively. The effects on the visual quality are shown in Figs. 7 (a)–(d), respectively.
[…]
((a))
((b))
Figure 8: Model compression performance.
@@ IV-C2 IMPACT OF MODEL COMPRESSION
IV-C3 IMPACT OF NETWORK ARCHITECTURE
Figure 9: Chamfer distance under the different patch sizes.
Refer to caption ((a)) Refer to caption ((b))
Refer to caption ((c)) Refer to caption ((d))
Refer to caption ((e)) Refer to caption ((f))
Refer to caption ((g))
Figure 10: Snapshots of the reconstructed LiDAR point clouds in proposed methods under the different patch sizes $N_{p}\times N_{p}$. Here, (b)-(g) show the reconstructed point clouds of frame 00 of sequence 00.
Figure 11: Chamfer distance under the different layer sizes.
Refer to caption ((a)) Refer to caption ((b))
Refer to caption ((c)) Refer to caption ((d))
Refer to caption ((e))
Figure 12: Snapshots of the reconstructed LiDAR point clouds in proposed methods under the different layer sizes $L$. Here, (b)-(e) show the reconstructed point clouds of frame 00 of sequence 00.
This section discusses the effect of the configurations for the depth INR architecture, specifically the patch size $N_{p}$ and layer size $L$, on the quality of the reconstructed LiDAR point cloud. The proposed depth INR uses the patch-wise architecture, and thus the depth image is divided into patches of size $N_{p}\times N_{p}$. Here, a small patch size increases the complexity of intra-patch learning, while a large patch size increases the complexity of inter-patch learning.
Fig. 9 shows CD performance varying $N_{p}$, and Fig. 10 demonstrates the corresponding snapshots of the reconstructed point cloud. We consider all the variants using the same model size in Fig. 8 with a sparsity of 0.0 and $N_{b}$ of 32. The evaluation results demonstrated that the patch size of $N_{p}=16$ yields the best CD performance. However, using larger or smaller $N_{p}$ values leads to performance degradation. While a large $N_{p}$ reduces the effectiveness of patch-wise modeling due to the limited pixel count, a small $N_{p}$ requires covering a wider area. This makes it challenging to capture sharp depth transitions and preserve geometric details near boundaries.
Similarly, Fig. 11 and 12 show the 3D reconstruction quality of the proposed scheme and the corresponding snapshots for different layer sizes $L$. The results indicate that a layer size of $L=6$ is the most effective for CD performance.
@@ V Conclusion and Future Work
We proposed a novel RI-based LiDAR point cloud compression method. The proposed method is designed to efficiently compress floating-point RIs using INR-based techniques and features a sophisticated architecture that combines separated learning for mask and depth images, patch-wise learning for depth images, and model compression. Experiments on the KITTI dataset show that the proposed method improves 3D reconstruction quality at low bitrates compared with conventional image codecs and representative baselines such as G-PCC, Draco, and COIN, and it also achieves strong 3D object detection accuracy in the low-bpp regime.
@@
The proposed method has two limitations: encoding delay and transformation loss from RIs to 3D point clouds. While existing baselines require only a few milliseconds for encoding, implicit neural compression, including the proposed method, takes from tens of minutes to several hours. In summary, the proposed method’s long encoding delay and short decoding delay make it well-suited for offline applications of LiDAR point clouds. To further reduce encoding latency, recent findings on learned initializations for coordinate-based neural representations [tancik2021learned] and meta-learned sparse INRs [lee2021meta] can be integrated into our depth/mask INR architecture. We leave the implementation and evaluation of such integration as future work.
In addition, RIs with limited spatial resolution will lead to irreversible point loss during 2D-to-3D decoding, potentially degrading the performance of downstream tasks. In future work, we will consider integrating a point cloud generator [2024arXiv240406715K] to obtain a denser point cloud from the limited resolution of RIs.
APPENDIX
This appendix provides further details for Table I. Table IV shows the detailed BD-CD performance across the different LiDAR frames.
TABLE IV: The list of BD-CD $\uparrow$ for the KITTI dataset across the different frames. Note that BD-CD is evaluated for each baseline using the proposed method as the reference. Positive values indicate that the proposed method achieves a lower chamfer distance than the corresponding baseline.
[…]
Oct
Attention
$\S$
R-PCC
$\P$
R-PCC
$\P$
00 00 1.197 0.661 0.465 0.260 1.039 0.099 0.146 -0.022 0.023 0.108
25 1.277 0.719 0.490 0.237 0.969 0.071 0.146 -0.032 0.028 0.118
50 1.588 0.923 0.607 0.266 1.264 0.081 0.153 -0.025 0.050 0.136
75 1.696 1.036 0.661 0.270 0.983 0.068 0.153 -0.021 0.016 0.100
100 1.208 0.667 0.450 0.247 1.261 0.088 0.189 -0.027 0.000 0.085
01 00 1.118 0.630 0.474 0.258 1.072 0.084 0.265 -0.036 -0.005 0.074
25 1.040 0.581 0.462 0.285 1.073 0.091 0.292 -0.013 0.005 0.082
50 1.128 0.624 0.485 0.271 1.113 0.121 0.238 -0.011 0.019 0.088
75 1.411 0.754 0.546 0.244 0.969 0.115 0.177 -0.005 0.021 0.080
100 1.472 0.842 0.606 0.239 1.250 0.075 0.199 -0.018 0.057 0.159
02 00 1.248 0.704 0.497 0.229 0.970 0.082 0.217 -0.056 -0.014 0.137
25 1.037 0.561 0.419 0.250 0.986 0.063 0.215 -0.031 0.002 0.070
50 1.039 0.558 0.418 0.266 0.996 0.099 0.272 -0.022 0.006 0.076
75 1.101 0.607 0.458 0.273 0.968 0.075 0.203 -0.017 0.009 0.087
100 1.176 0.629 0.446 0.250 1.276 0.095 0.223 -0.005 0.027 0.101
03 00 1.524 0.847 0.551 0.228 1.037 0.062 0.164 -0.031 0.036 0.124
25 1.559 0.973 0.615 0.244 0.960 0.061 0.139 -0.040 0.019 0.109
50 1.293 0.737 0.496 0.222 1.215 0.050 0.203 -0.055 -0.019 0.071
75 1.129 0.560 0.405 0.232 0.927 0.060 0.202 -0.055 -0.041 0.036
100 1.103 0.548 0.421 0.237 1.232 0.063 0.174 -0.045 -0.029 0.052
04 00 1.078 0.566 0.420 0.223 1.046 0.043 0.210 -0.072 -0.002 0.115
25 1.348 0.806 0.546 0.239 1.045 0.085 0.199 -0.052 -0.012 0.065
50 1.594 0.876 0.607 0.241 1.131 0.090 0.196 -0.020 0.028 0.105
75 1.391 0.789 0.518 0.230 0.951 0.062 0.154 -0.044 0.005 0.089
100 1.307 0.704 0.483 0.252 1.288 0.088 0.199 -0.005 0.030 0.109
05 00 1.108 0.620 0.439 0.252 1.007 0.083 0.179 -0.038 0.032 0.135
25 1.198 0.661 0.464 0.256 1.009 0.077 0.187 -0.027 -0.005 0.071
50 1.774 1.078 0.679 0.259 0.997 0.059 0.125 -0.043 0.002 0.084
75 1.752 1.100 0.698 0.241 0.935 0.051 0.109 -0.058 -0.002 0.084
100 1.589 0.924 0.601 0.233 1.258 0.055 0.166 -0.036 0.012 0.102
06 00 1.161 0.686 0.471 0.224 0.954 0.052 0.123 -0.069 0.010 0.122
25 1.439 0.824 0.530 0.218 0.950 0.048 0.184 -0.068 -0.027 0.059
50 1.359 0.814 0.531 0.226 0.951 0.046 0.159 -0.070 -0.031 0.045
75 1.532 0.976 0.654 0.223 0.905 0.038 0.161 -0.079 -0.028 0.047
100 1.128 0.623 0.451 0.230 1.205 0.082 0.141 -0.061 -0.045 0.035
Average 1.317 0.749 0.516 0.244 1.063 0.073 0.185 -0.037 0.005 0.090
  • $\dagger$: Image based method(s),   $\ddagger$}: INR based method,   $\S$: Point Cloud based method(s),   $\P$: Range Image based method(s).
Acknowledgment
This work was supported by JST·ASPIRE Grant Number JPMJAP2432 and JSPS KAKENHI Grant Number JP22H03582.